solve each equation
1) x^2+x-12=0
2) 2a^2-8a=0

Answer by benni1013(206)

solve each equation
1) x^2+x-12=0
2) 2a^2-8a=0
I will do one for you and you do the other.
1) x^2+x-12=0. Ask yourself the question of what factors of -12 equals +x
Ans: 4,-3
x^2+4x-3x-12=0==>Take out common factor of x==>x(x+4) and take out common factor of -3 -3(x+4)==>x(x+4)-3(x+4)=0==>Since we have common terms in the parentheses we can bring everything together==>(x-3)(x+4)=0==>Divide and conquer
x-3=0 and x+4=0==>Solve==>x=3,4